OpenStudy (anonymous):

If f(x) = x^2 − x + 6, what is the value of f(x + h) − f(x)/h? x + 6 x − 1 2x − 1 2x + 6 x + 5

Parth (parthkohli):

What did you try?

OpenStudy (anonymous):

i dont know this kind of stuff so i havent tried

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

It must have been taught to you during the lessons. I suggest you to look over them and at least try something.

OpenStudy (anonymous):

im getting ahead in school so its online

OpenStudy (skullpatrol):

what does this expression f(x + h) − f(x)/h mean to you?

OpenStudy (anonymous):

i really dont know?

OpenStudy (skullpatrol):

$$\Huge \frac{f(x + h) − f(x)}{h}$$

OpenStudy (anonymous):

i understand this but nothing else

OpenStudy (skullpatrol):

$$\Huge f(x) = x^2 − x + 6$$ $$\Huge f(x+h) = (x+h)^2 −( x+h ) + 6$$

OpenStudy (skullpatrol):

Does that^ make sense?

OpenStudy (anonymous):

yeah

OpenStudy (skullpatrol):

Ok, multiply it out and substitute in f(x + h) − f(x)/h and simply :-)

OpenStudy (anonymous):

i got 2x+6 but i dont know if its right

OpenStudy (skullpatrol):